Circular No. 7124

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999an IN IC 755 J. Y. Wei, L. Cao, Y. L. Qiu, Q. Y. Qiao, and J. Y. Hu, Beijing Astro-

nomical Observatory (BAO), report the discovery by the BAO Supernova Survey of a supernova (unfiltered mag ∼ 15) with the 0.6-m telescope at Xinglong Station on Mar. 7.83 UT. SN 1999an is located at α = 12^{h}01^{m}10^{s}.5, δ = +14^{o}06′11′′ (equinox 2000.0), close to the nucleus of IC 755. SN 1999an also appears on images taken on Jan. 22.85, 26.84, Feb. 7.89, and 15.79. A low-resolution spectrum (range 400–820 nm), obtained by L. Cao and Q. S. Gu with the BAO 2.16-m telescope on Mar. 10.8, shows SN 1999an to be a type-II supernova with obvious P-Cyg Balmer lines. Strong, narrow emission lines from H II regions also appear in the spectrum and show that IC 755 has a redshift of 0.004. A photospheric expansion velocity of 5300 km/s was obtained from the Hα P-Cyg profile.

SUPERNOVA 1999ao IN ANONYMOUS GALAXY D. Reiss, University of Washington, and the Mt. Stromlo Abell Cluster

Supernova Search Team (cf. IAUC 7106) report the discovery of a supernova (V ∼ 18.2, R ∼ 18.4) on a CCD image obtained on Mar. 3 by S. Sabine with the Mt. Stromlo 1.27-m telescope (+ Macho Camera). SN 1999ao is located near Abell 3392 at α = 6^{h}27^{m}26^{s}.37, δ = −35^{o}50′24′′.2 (equinox 2000.0), which is 2′′.3 west and 0′′.9 north of the nucleus of the host galaxy; a nearby star has position end figures 27^{s}.34, 51′09′′.3. SN 1999ao is not visible on similar frames taken on Jan. 10 (limiting mag 20.5). CCD UBVRI images obtained at the Cerro Tololo 0.9-m telescope on Mar. 8 and 9 by L. Germany also show the new star. A spectrum (range 320–900 nm) taken by B. Schmidt at the Australian National University’s 2.3-m telescope on Mar. 10.5 UT shows that SN 1999ao is a type-Ia event, ∼ 8 days before maximum brightness; its redshift (z = 0.054), obtained from Ca H and K features, is nearly identical to that of Abell 3392 (z = 0.055).

NOVA MUSCAE 1998 A. Retter, Keele University; W. Liller, Viña Del Mar, Chile; and G.

Garradd, Loomberah, N.S.W., write: “Analysis of CCD continuous photometry of N Mus 1998 through wide-band filters during 15 nights in Jan.– Feb. reveals a periodic oscillation. The period is either 0.16930 ± 0.00015 day or its 1-day alias of 0.20380 ± 0.00025 day. The semiamplitude of the variation is ∼ 0.03 mag. Most of the observations were carried out during the transition phase of the nova.

1999 March 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green